# Three-dimensional thermodynamic structures of the intracluster medium across edges in the X-ray surface brightness of massive, bright, dynamically-active galaxy clusters

### v1

@@ 3 Observations and data reduction
We use the archival X-ray datasets of our sample taken with the Advanced CCD Imaging Spectrometer (ACIS; Garmire et al., 2003) on board the Chandra X-ray Observatory. All datasets analyzed are summarized in Table $1$In this paper, we do not use datasets taken from deep 2 Msec Chandra observations of A2146 conducted in 2018 and 2019. Russell et al. (2022) investigated systematic uncertainty arising from the stability of the focal-plane temperature of the ACIS for these 2 Msec datasets. They found that the ICM temperature and abundance are consistent with those derived from the datasets of earlier observations, but the ICM electron density became $\sim 2$ higher. Since our forward modeling appraoch simultaneously estimates both the ICM density and temperature, this level of the systematic uncertainty impacts on our analysis. Therefore, we use only the first $\sim 420$ ksec datasets in this paper.. We use the versions of 4.15 and 4.10.2 for Chandra Interactive Analysis of Observations (CIAO; Fruscione et al., 2006) and the calibration database (CALDB), respectively. After applying the lc_clean task in CIAO to the data to exclude the duration of the flare, we obtain the net exposure times for each cluster summarized in Table 1. We adopt the blank-sky data included in the CALDB to estimate the background contribution to each galaxy cluster. We extract the X-ray spectrum of the ICM from the region of interest in each dataset with the specexctract task in CIAO and combined them after making individual spectrum, response, and ancillary response files for the spectral fitting. We use XSPEC version 12.13.0c (Arnaud, 1996) and the atomic database for plasma emission modeling version 3.0.9 in the X-ray spectral analysis, assuming that the ICM is in collisional ionization equilibration. Thus, we use the APEC model in XSPEC (Smith et al., 2001; Foster et al., 2012). The abundance table of Anders & Grevesse (1989) is used. Here, the abundance of a given element is defined as $Z_{i}=(n_{i,{\rm obs}}/n_{\rm H,obs})/(n_{i,\odot}/n_{\rm H,\odot})$, where $n_{i}$ and $n_{\rm H}$ are the number densities of the $i$th element and hydrogen, respectively. We use the iron abundance to represent the ICM metal abundance, such that the abundance of other elements is tied to the iron abundance as $Z_{i}=Z_{\rm Fe}$ (Ueda et al., 2021). The Galactic absorption (i.e., $N_{\mathrm{H}}$) for each galaxy cluster is estimated using HI4PI Collaboration et al. (2016) and fixed at the estimated value in the X-ray analysis as follows: $N_{\mathrm{H}}$ $=4.25\times 10^{20}$ cm-2 for A3667, $N_{\mathrm{H}}$ $=1.36\times 10^{21}$ cm-2 for A2319, $N_{\mathrm{H}}$ $=5.51\times 10^{20}$ cm-2 for A520, and $N_{\mathrm{H}}$ $=2.62\times 10^{20}$ cm-2 for A2146.
@@ 4 Analyses and Results
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 4.1 X-ray images of our sample
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 5.3 Azimuthal distributions of the ICM thermal pressure ratio at the cold fronts
The higher pressure jump at the interface in the $180^{\circ}-210^{\circ}$ sector of A2319 may be a hint of exploring the mechanisms to create cold fronts. If gas sloshing is the case, this sector is located at the head of the spiral pattern (Ichinohe et al., 2021). The head of a sloshing spiral is expected to have the fastest tangential gas flow. In fact, transonic motions associated with gas sloshing is found in Abell 907 (Ueda et al., 2019). In this context, the higher pressure jump may be evidence of gas sloshing. On the other hand, although the cold front in A3667 is widely recognized as a stripping cold front, the hypothesis that this cold front results from gas sloshing either in the plane of the sky or the LOS can be considered. Since no apparent difference in the trend for the pressure ratio is found, it is difficult to distinguish stripping cold fronts from sloshing cold fronts, based on the results obtained from the forward modeling analysis. This hypothesis of gas sloshing is testable with XRISM.
@@ 5.4 Position of the interface
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 6 Summary and conclusions
We thank Maxim Markevitch for fruitful discussions and constructive comments. We also thank Keiichi Umetsu and Sandor Molnar for their helpful discussions. The scientific results reported in this article are based on data obtained from the Chandra Data Archive. SU acknowledges the support from the National Science and Technology Council of Taiwan (NSTC 111-2112-M-001-026-MY3).
@@ Appendix B Radial profiles
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption

### v2

@@ 3 Observations and data reduction
We use the archival X-ray datasets of our sample taken with the Advanced CCD Imaging Spectrometer (ACIS; Garmire et al., 2003; Grant et al., 2024) on board the Chandra X-ray Observatory. This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in https://doi.org/10.25574/cdc.265 (catalog DOI: 10.25574/cdc.265). All datasets analyzed are summarized in Table $1$In this paper, we do not use datasets taken from deep 2 Msec Chandra observations of A2146 conducted in 2018 and 2019. Russell et al. (2022) investigated systematic uncertainty arising from the stability of the focal-plane temperature of the ACIS for these 2 Msec datasets. They found that the ICM temperature and abundance are consistent with those derived from the datasets of earlier observations, but the ICM electron density became $\sim 2$ higher. Since our forward modeling appraoch simultaneously estimates both the ICM density and temperature, this level of the systematic uncertainty impacts on our analysis. Therefore, we use only the first $\sim 420$ ksec datasets in this paper.. We use the versions of 4.15 and 4.10.2 for Chandra Interactive Analysis of Observations (CIAO; Fruscione et al., 2006) and the calibration database (CALDB), respectively. After applying the lc_clean task in CIAO to the data to exclude the duration of the flare, we obtain the net exposure times for each cluster summarized in Table 1. We adopt the blank-sky data included in the CALDB to estimate the background contribution to each galaxy cluster. We extract the X-ray spectrum of the ICM from the region of interest in each dataset with the specexctract task in CIAO and combined them after making individual spectrum, response, and ancillary response files for the spectral fitting. We use XSPEC version 12.13.0c (Arnaud, 1996) and the atomic database for plasma emission modeling version 3.0.9 in the X-ray spectral analysis, assuming that the ICM is in collisional ionization equilibration. Thus, we use the APEC model in XSPEC (Smith et al., 2001; Foster et al., 2012). The abundance table of Anders & Grevesse (1989) is used. Here, the abundance of a given element is defined as $Z_{i}=(n_{i,{\rm obs}}/n_{\rm H,obs})/(n_{i,\odot}/n_{\rm H,\odot})$, where $n_{i}$ and $n_{\rm H}$ are the number densities of the $i$th element and hydrogen, respectively. We use the iron abundance to represent the ICM metal abundance, such that the abundance of other elements is tied to the iron abundance as $Z_{i}=Z_{\rm Fe}$ (Ueda et al., 2021). The Galactic absorption (i.e., $N_{\mathrm{H}}$) for each galaxy cluster is estimated using HI4PI Collaboration et al. (2016) and fixed at the estimated value in the X-ray analysis as follows: $N_{\mathrm{H}}$ $=4.25\times 10^{20}$ cm-2 for A3667, $N_{\mathrm{H}}$ $=1.36\times 10^{21}$ cm-2 for A2319, $N_{\mathrm{H}}$ $=5.51\times 10^{20}$ cm-2 for A520, and $N_{\mathrm{H}}$ $=2.62\times 10^{20}$ cm-2 for A2146.
@@ 4.2 Parametric forward fitting
We generate exposure maps of each energy band using the fluximage task in CIAO. We set an effective energy for each exposure map to the midpoint of the corresponding energy band. Thus, we extract the radial profiles of the exposure maps using the same approach as the X-ray surface brightness.
@@ 5 Discussion
In this paper, we have assumed an isothermal temperature profile for the outer regions, i.e., regions above the interface (see Equation 2), to reduce computational time. However, assuming such an isothermal temperature profile may be simplistic. In general, the ICM temperature decreases toward the outskirts. To investigate the systematic uncertainty due to the assumption of an isothermal temperature profile, we analyze the sectors exhibiting the highest Mach number in A520 and A2146, respectively, with a power-law temperature profile for the outer region (i.e., $T(r)=T_{\rm out}(r/r_{23})^{\gamma}:r_{23}<r$), instead of the isothermal temperature profile. We find no significant difference between the results derived from the isothermal and power-law temperature profiles. Further study is needed to extend a power-law temperature profile for our forward modeling approach and apply it to other interfaces.
@@ 5.3 Azimuthal distributions of the ICM thermal pressure ratio at the cold fronts
The higher pressure jump at the interface in the $180^{\circ}-210^{\circ}$ sector of A2319 may be a hint of exploring the mechanisms to create cold fronts. If gas sloshing is the case, this sector is located at the head of the spiral pattern (Ichinohe et al., 2021). The head of a sloshing spiral is expected to have the fastest tangential gas flow. In fact, transonic motions associated with gas sloshing are found in Abell 907 (Ueda et al., 2019). In this context, the higher pressure jump may be evidence of gas sloshing. However, the presence of clear X-ray substructures in the north-west direction may indicate a major merger in A2319 (e.g., Markevitch, 1996; Ghirardini et al., 2018). In fact, the observed flat entropy profiles in the inner region of A2319 support this hypothesis. If a major merger is the case, the spiral pattern can be interpreted as bulk motions of the core of A2319, in a similar fashion to those of A754 (e.g., Markevitch et al., 2003; Henry et al., 2004; Macario et al., 2011). Further study is needed to explore the origin of the spiral pattern. On the other hand, although the cold front in A3667 is widely recognized as a stripping cold front, the hypothesis that this cold front results from gas sloshing either in the plane of the sky or the LOS can be considered. Since no apparent difference in the trend for the pressure ratio is found, it is difficult to distinguish stripping cold fronts from sloshing cold fronts, based on the results obtained from the forward modeling analysis. This hypothesis of gas sloshing is testable with XRISM.
@@ 6 Summary and conclusions
We are grateful to the anonymous referee for the helpful and constructive comments. We thank Maxim Markevitch for the fruitful discussions and constructive comments. We also thank Keiichi Umetsu and Sandor Molnar for their helpful discussions. The scientific results reported in this article are based on data obtained from the Chandra Data Archive [^2] https://doi.org/10.25574/cdc.265. SU acknowledges the support from the National Science and Technology Council of Taiwan (NSTC 111-2112-M-001-026-MY3).